OpenStudy (jjuden):

Evaluate 3+sqrt of x+7 - sqrt 3x when x = 9. Write the answer as a decimal. Round to the nearest hundredth if necessary.

OpenStudy (johnweldon1993):

\[\large 3 + \sqrt{x+7} - \sqrt{3x}\] Like that?

OpenStudy (jjuden):

yes

OpenStudy (anonymous):

so... sub in your 9 in replace of all x's and show us the new equation!

OpenStudy (johnweldon1993):

Okay...so when x = 9 you plug in 9 for every 'x' in that equation \[\large 3 + \sqrt{9 + 7} - \sqrt{3(9)}\] \[\large 3 + \sqrt{16} - \sqrt{27}\] \[\large 3 + 4 - \sqrt{27}\] Can you solve it from here?

OpenStudy (jjuden):

no i'm terrible at this